Find the self-inductance of a solenoid 70 cm long and 5.8 cm in diameter that contains 2000 turns of wire.
Answer #1

L = uN^2A/l

= 4*3.14*10^-7 * 2000^2 * 3.14*(.058/2)^2/.7

= .0189 H = 18.9 *10^-3 H
Answer #2

L = (N^2)(mu)(Area)/(length)
mu = 4? x 10^-7 henry per meter
L = (2000)^2(4? x 10^-7)(5.8)/(70)
L = 0.4163 henry
Answer #3

media%2Fc03%2Fc0379904-26b3-482b-b34c-27

The self induction of the solenoid is μ.N2 A (47x10 T.mA)(2000). (2.640 x 10 m*) 70x10 m -1895 x10 H

Answer question